Elementary Functions and Calculus II Math 132 (Sec 50), Winter 2005 Course Document

Lecturer: David Rule Office: Eckhart 16 Phone: (773) 702 6045 Email: rule@math.uchicago.edu Website: www.math.uchicago.edu/˜ rule/132/

Lectures: Monday, Wednesday & Friday, 12:30am-1:20pm, Room SS105.

Tutorials: Tuesday & Thursday, 4:30pm-5:50pm David Lawlor (dlawlor@uchicago.edu) C116 Sara Revzi (arsinoe@uchicago.edu) C106 Louis Sloniker (sloniker@uchicago.edu) C103

Office hours: TBA and by appointment

Text: Calculus, Varberg, Purcell & Rigdon, Chapters 4, 5, 6 & 7.

Syllabus: The 131-2-3 sequence aims to teach the basic concepts and techniques of rigorous mathematics and apply these to certain problems. Included in these aims is presenting you work in an understandable and coherent manner. The course 132 will cover trigonometric functions and other transcendental functions, the mean value theorem for derivatives, primitives, the Riemann integral, the mean value theorem for integrals, the fundamental theorems of calculus and applications of the integral.

Assessment: Your grade will be determined by two in-class mid-term tests, a final examination, and homework. I may occasionally give quizzes but these will not count towards your grade. Homework should be handed in on or before the given deadline. There will be no extensions given for homework, however, the worst of your homework scores will not count towards your final homework score. The final constitutes 40% of your grade, the midterms 20% each, and homework totals 20%.

The Final Examination: The final examination is on Thursday, 17th March from 4pm to 6pm. It is the policy of the Department of Mathematics that the following rules apply to final exams in all undergraduate mathematics courses:

1. The final exam must occur at the time and place designated on the College Final Exam Schedule. In particular, no final examinations may be given during the tenth week of the quarter, except in the case of graduating seniors.

2. Any student who wishes to depart from the scheduled final exam time for the course must receive permission from Paul Sally (office is Ry 350, phone is 2-7388, email is sally@math.uchicago.edu). Instructors are not permitted to excuse students from the scheduled time of the final exam except in the cases of an Incomplete. The University examination schedule can be found at http://registrar.uchicago.edu/students/final-exams.html.

Miscellaneous Items: The use of calculators, both for homework and during examinations, is forbidden. The rule is a little redundant as you will find them next to useless.

David Rule